# Temporal Feature Matters: A Framework for Diffusion Model Quantization

### v3

@@ 5.3.1 Efficiency of TM
First, compared with BRECQ, our TIB-based Maintenance brings less than 0.08% extra memory consumption, since we employ per-tensor quantization with our FSC for activations. Besides, our strategy only induces an approximated 0.3% speed overhead for Stable Diffusion and SD-XL on CPUs, and the latency comparison results on the GPU also show nearly no difference.
5.3.2 Efficiency of CM
@@
Second, our Cache-based Maintenance removes the necessity to compute temporal features online, which slightly accelerates inference. Specifically, 4.33% speed enhancement for SD-XL on CPU is obtained compared with TM. Note that, CM also requires less storage here, e.g., 26.65MB memory reduction for Stable Diffusion than TM. Even for some models with a large time-steps count like 500 steps LDM on LSUN-Bedroom $256\times 256$, an additional 10.4MB memory requirement without the TIB brings an extra cost amounting to roughly 3.69% of the quantized UNet size (281.8MB). Moreover, our highly compatible framework can seamlessly integrate step-reduction techniques like advanced sampler or step distillation to further speed up and reduce memory consumption. We leave this for our future study.
5.3.3 Efficiency of DS
@@
Last, equipping with the Disturbance-aware Selection, our complete framework achieves a pronounced efficient inference with significantly lower memory consumption compared with FP. For example, the quantized SD-XL is 5.76$\times$ faster than the full precision with 3.99$\times$ memory savings. It is obvious that the time consumption of all different selection results is between TM and CM, so less than a 5% gap between these results is achieved as shown in Fig. 10. Besides, compared to TM in Fig. 9, DS has almost no additional overhead. Based on these analyses, we verify the high efficiency of our framework for inference and deployment.
5.4 Ablation Studies
To evaluate the effectiveness of different modules in our proposed method, we perform a thorough ablation study on the LSUN-Bedrooms $256\times 256$ dataset with w4a8 quantization, utilizing the LDM-4 model with a DDIM sampler, unless otherwise stated. The basic ablation is shown in Tab. VI.
@@ 5.4 Ablation Studies
The content is organized as follows. We first explore the effect of TIB-based Maintenance in Sec. 5.4.1. Then, we present detailed calibration ablation for TIB-based Maintenance and Cache-based Maintenance in Sec. 5.4.2 and Sec. 5.4.3, respectively. Moreover, we evaluate the loss function of Disturbance-aware Selection and detailed analyses of its selection proportion in Sec. 5.4.4 and Sec. 5.4.5. Finally, the performance with advanced samplers is exhibited to further showcase our universality in Sec. 5.4.6.
5.4.1 Effect of TM
As shown in Tab. VI, compared to the baseline method, our TIAR reduces FID and sFID by 4.52 and 13.44, respectively. Additionally, our FSC also achieves promising results. Equipping both approaches can further mitigate performance degradation.
@@ 5.4.1 Effect of TM
For a more detailed analysis, Fig. 11 (Left) reveals that our methods induce significantly less temporal feature disturbance compared to PTQD. This underscores the efficacy of our motivation and contribution to maintaining temporal features. Further insights are gained by examining the cosine similarity between outputs from the $i^{\text{th}}$ Residual Bottleneck Blocks pre- and post-quantization. Comparing our methods with PTQD, Fig. 11 (Right) illustrates that our approaches result in significantly lower output errors in the Residual Bottleneck Block. It is also essential to note that these points involve the accumulated errors from multiple denoising iterations in diffusion models. Therefore, Our methods are able to significantly eliminate the accumulative error and quantization error by mitigating temporal feature disturbance within the corresponding block for performance enhancements.
5.4.2 Effect of Calibration Methods for TM
Given the variety of available methods to determine the optimal step size for activations during calibration, we explore several approaches and evaluate both their performance and the GPU time consumed. As detailed in Tab. VII, the performance differences between Min-max and the best-performing ones are minimal, e.g., a mere 0.06 FID gap under w4a8. Therefore, we opt for the simplest and most efficient Min-max [62] as our specific calibration strategy, striking a balance between calibration time and effectiveness. Notably, we have found that PTQD and Q-Diffusion cost 4.68 and 5.29 GPU hours in their PTQ methods under w4a8 quantization, respectively. However, our TIB-based maintenance only spends 2.32 GPU hours (2$\times$ speedup) due to our efficient calibration and reconstruction without extra time-consuming compared with previous block/layer-wise reconstruction.
TABLE VII: Different range estimation methods for calibration of TIB-based Maintenance. We report the GPU hours consumed during calibration in the table. Weight reconstruction combined with our TIAR (2.20 GPU hours) is employed before calibration combined with FSC. The subscript numbers represent the improvements between the best-performing methods and Min-max [62].
@@ 5.4.2 Effect of Calibration Methods for TM
5.4.3 Effect of Calibration Methods for CM
@@
We also investigate different calibration methods for our Cache-based Maintenance. From Tab. VIII, we observe that LSQ outperforms the others, achieving a 13.3% increase in SQNR compared to MSE under w8a8 settings. Hence we choose LSQ as our calibration method. Note that each pre-computed temporal feature can be processed in parallel, enabling the algorithm to achieve rapid runtimes of less than 0.3 hours on a single H800 80G GPU.
5.4.4 Effect of Loss Functions for DS
Since the loss function $\mathcal{L}_{t,i}^{temporal}$ in Eq. (14) measures the error of the temporal feature, we assess various loss functions to enhance performance. As shown in Tab. IX, different loss functions produce comparable results, with less than 1% FID and sFID discrepancies among them. Moreover, all results in the table outperform either TIB-based or Cache-based Maintenance in Tab. II, demonstrating the robustness and effectiveness of the selection strategy.
TABLE IX: Effect of different loss functions in Eq. (14). We employ MSE [104] align with Eq. (13) in our framework.
@@ 5.4.4 Effect of Loss Functions for DS
5.4.5 Effect of Selection Proportion
Beginning with Cache-based Maintenance, we select a proportion of the temporal feature with the lowest $\tau_{t,i}$ for TIB-based Maintenance. Then we increase the proportion progressively until the selection pattern is equivalent to TIB-based Maintenance. As shown in Fig. 12, the lowest FID and highest SQNR achieved by DS compared with other points. Furthermore, selecting maintenance strategies that minimize temporal feature errors can help enhance performance. This investigation also validates our fundamental perspective that temporal feature maintenance is crucial.
Figure 12: Performance from Cache-based Maintenance $\rightarrow$ Disturbance-aware Selection $\rightarrow$ TIB-based Maintenance. The x-axis represents the proportion of temporal features applied using TIB-based Maintenance.
5.4.6 Generation with Advanced Samplers
Apart from using the DDIM sampler [82], we also utilize a variant of DDPM [22] called PLMS [51] on the CelebA-HQ $256\times 256$ dataset [30]. This better demonstrates the superiority of our framework compared to previous works. From Tab. X, the introduced DS substantially reduces FID and sFID, surpassing PTQD by margins of 12.61 and 7.08, respectively.
@@ 5.4.6 Generation with Advanced Samplers
Furthermore, we present experiments performed using the DPM++ solver [55] on LSUN-Churches $256\times 256$ [95]. As shown in Tab. XI, our framework consistently outperforms existing methods.
6 Conclusion & Discussions of Limitations
@@
However, beyond temporal features, we have discovered that other semantic features introduced in conditional diffusion models possess physical significance and impact the generation effect. Nevertheless, these features are often overlooked in current methodologies. We plan to address this in future research. Furthermore, while temporal feature maintenance is effective in both PTQ and QAT scenarios, our current study focuses primarily on the PTQ setting. Future efforts will extend to the QAT setting to pursue lower-bit quantization and further performance improvements.
Supplementary Material for “Temporal Feature Matters: A Framework for Diffusion Model Quantization”
We organize the supplementary material as follows.
  • In Sec. A, we provide a detailed relationship between temporal feature errors and final generation quality.
  • In Sec. B, we conduct more experiments to validate the inappropriate reconstruction target of previous methods.
  • In Sec. C, we explore the effectiveness of our framework on video generation.
  • In Sec. D, we provide a guideline on when to use which proposed strategies.
  • In Sec. E, we conduct experiments to show the superiority of our framework compared to previous PTQ methods under sub-4-bit quantization.
  • In Sec. F, we apply our framework to the advanced step-distilled model (i.e., FLUX.1-Schnell).
  • In Sec. G, we deploy our quantized models on edge and mobile devices.
@@ A Relationship between Temporal Feature Errors and Final Generation Quality
In this section, we sort out the relationship between temporal feature error and final generation quality. First, we define temporal feature disturbance as a significant temporal feature error (Sec. 4.1). Then we show that temporal feature disturbance (i.e., significant temporal feature error) accounts for temporal information mismatch (Sec. 4.2), which leads to trajectory deviation (Sec. 4.2). In Fig. 4, the deviated trajectory can result in noisy images gradually shifting from a normal denoising trajectory. Thus, the image quality degrades severely. Moreover, as reflected in Fig. 2, with temporal feature error increasing [^13] Obviously, temporal feature error which can also be defined by $\mathcal{L}^{temporal}_{t,i}$ in Eq. (13) increases with the increase of noise level $\lambda$., generation quality measured by FID score deteriorates sharper and sharper.
@@ B Inappropriate Reconstruction Target Validation
We further conduct the validation experiments for the baselines (i.e., PTQ4DM [79], Q-Diffusion [44], and PTQD [19]) in Tab. II. As shown in Tab. I, the Freeze strategy still significantly improves the performance akin to that in Tab. I. For example, it decreases FID and sFID by 1.18 and 5.92 for w4a8 Q-Diffusion. Hence, the results further validate our analysis and make it more convincing.
@@ C Exploration for Video Generation
Advanced Video diffusion models employ DiT [69] as their backbones. We have found that this type of backbone has temporal features that can be formulated by Eq. (6), where $g_{i}(\cdot)$ is a simple multi-layer perception (MLP) network. Moreover, temporal features and modules generating them are also independent of any $\mathbf{x}_{t}$. Thus, we can directly use our framework for these video diffusion models. Due to similar architectures and algorithms, we believe our approach can also achieve superior performance. As shown in Tab. II, we conduct experiments on OpenSora [103] with 100-steps DDIM [82] and cfg scale of 4.0. We select representative metrics and measure them on 10 examples in OpenSora prompts set [103] the same as the previous study [100]. Specifically, following EvalCrafter [52], we select CLIPSIM and CLIP-Temp to measure the text-video alignment and temporal semantic consistency. DOVER [92]’s video quality assessment (VQA) metrics to evaluate the generation quality from aesthetic and technical aspects are also involved. $\Delta$Flow Score [52] is used for evaluating the temporal consistency. We employ 8 prompts from the OpenSora prompt set to quantize the model. Other quantization settings are the same as that of PTQ4DiT [93]. Compared with methods tailored for DiT architecture in Tab. II, our framework outperforms them across all metrics. We will explore more about video generation in the future.
@@ D Guidelines on When to Use Which Proposed Strategies
We provide a guideline here on when to use which maintenance strategy. First, we clarify that the time overhead  [^14] The additional time overhead only related to the process to obtain quantized temporal features is unrelated to $T$. of the two maintenance strategies is all less than 5% as shown in Fig. 10. Only considering memory overhead, both methods might be suitable for different scenarios. For example, to accommodate a dynamic timestep range of 1$\sim$1000, solely storing temporal features without $\{g_{i}\}_{i=1,\ldots,n}\cup\{h\}$ (i.e., Cache-based Maintenance) for w4a8 LDM-8 on LSUN-Churches $256\times 256$ [95] requires an additional 20.7MB. It is a significant extra cost amounting to roughly 14.7% of the quantized UNet size (i.e., 140.9MB). In contrast, TIB-based Maintenance barely incurses less than 1% memory overhead in this scenario. However, Cache-based Maintenance can also require less storage in other cases, e.g., 26.65MB memory reduction for Stable-Diffusion than TIB-based Maintenance as depicted in Fig. 9. In conclusion, users can choose these two maintenance methods for deployment according to their specific hardware, diffusion backbone, and inference arguments etc. Without considering any resource restriction, we suggest users use both methods with our selection approach for higher performance.
@@ E Sub-4-bit Quantization Results
TABLE III: Sub-4-bit quantization results for unconditional/class-conditional image generation with LDM. Experimental details are the same as that in the main text.
@@
TM 2/4 38.44 -8.32 162.33 -43.00 174.42 -42.64 64.20 -14.24 29.05 -8.58 50.01 -5.59 35.44 -8.79
CM 2/4 42.37 -12.25 150.21 -55.12 179.04 -38.02 60.24 -18.20 28.41 -9.22 49.46 -6.14 33.90 -10.33
DS 2/4 46.02 -12.25 147.32 -58.01 170.11 -46.95 56.38 -22.06 27.24 -10.39 47.42 -8.18 31.04 -13.19
For sub-4-bit settings, we employ w4a4 and w2a4 quantization with LDM on ImageNet 256$\times$256 [10], LSUN-Bedrooms 256$\times$256 [10], and LSUN-Churches 256$\times$256 [95]. The results in Tab. III show the significant performance improvement achieved by our method compared with other PTQ baselines. Specifically, the superiority of our method to other baselines widens at the decrease of the bit-width. However, we have to clarify that most current hardware does not support sub-4-bit quantization. Moreover, in these extremely low-bit settings, QAT, which also re-trains or fine-turns model weights, is a more common choice. As discussed in the conclusion section, we believe the insight of temporal feature maintenance can also be generalized to QAT settings for further improving sub-4-bit quantization performance. This can be left as one of the future work.
@@
As shown in Tab. V, our method for SD-XL-turbo [78] with a single sampling step distilled from SD-XL-base [71] shows satisfactory boosts. The improvement is even $>$2 points larger than that for the 50-step SD-XL-base under the w4a8 setting. Here, We further conduct experiments for FLUX.1-Schnell [42] in this section. Specifically, we employ 512 COCO prompts for quantization. Other settings are the same as that of SD-XL-turbo [78]. The results in Tab. IV further validate the essential role of temporal features in this kind of scenario.
@@
Methods #Bits (W/A) 1-Step FLUX.1-Schnell 4-Step FLUX.1-Schnell

### v4

@@ V-C1 Efficiency of TM
First, compared with BRECQ, our TIB-based Maintenance brings less than $\mathbf{0.08\%}$ extra memory consumption, since we employ per-tensor quantization with our FSC for activations. Besides, our strategy only induces an approximated $\mathbf{0.3\%}$ speed overhead for Stable Diffusion and SD-XL on CPUs, and the latency comparison results on the GPU also show nearly no difference.
V-C2 Efficiency of CM
@@
Second, our CM removes the necessity to compute temporal features online, which slightly accelerates inference. Specifically, $\mathbf{4.33\%}$ speed enhancement for SD-XL on CPU is obtained compared with TM. Note that CM also requires less storage here, e.g., $\mathbf{26.65}$MB memory reduction for Stable Diffusion than TM. Even for some models with a large timesteps count, like $500$ steps for LDM on LSUN-Bedroom $256\times 256$, an additional $\mathbf{10.4}$MB memory requirement without the TIB brings an extra cost amounting to roughly $\mathbf{3.69\%}$ of the quantized UNet size ($281.8$MB). Moreover, our framework can seamlessly integrate step-reduction techniques like advanced samplers or step distillation to reduce inference costs. We leave this for our future study.
V-C3 Efficiency of DS
@@
Last, equipped with the DS, our complete framework achieves a pronounced efficient inference with significantly lower memory consumption compared with FP. For example, the quantized SD-XL is $\mathbf{5.76\times}$ faster than the full precision with $\mathbf{3.99\times}$ memory savings. It is obvious that the time consumption of all different selection results is between TM and CM, so less than a $\mathbf{5\%}$ gap between these results is achieved as shown in Fig. 9. Besides, compared to TM in Fig. 10, DS has almost no additional overhead. Thus, we verify the high inference efficiency of our framework.
V-D Ablation Studies
@@
To evaluate the effectiveness of different modules in our proposed method, we perform a thorough ablation study on the LSUN-Bedrooms $256\times 256$ dataset with W4444A8888 quantization, utilizing the LDM-4 model with a DDIM sampler, unless otherwise stated. The basic ablation is shown in Tab. VI.
The content is organized as follows. We first explore the effect of TIB-based Maintenance in Sec. V-D1. Then, we present detailed calibration ablation for TIB-based Maintenance and Cache-based Maintenance in Sec. V-D2 and Sec. V-D3, respectively. Moreover, we evaluate the loss function of Disturbance-aware Selection and conduct detailed analyses of its selection proportion in Sec. V-D4 and Sec. V-D5. Finally, the performance with advanced samplers is exhibited to further showcase our universality in Sec. V-D6.
V-D1 Effect of TM
As shown in Tab. VI, compared to the baseline method, our TIAR reduces FID and sFID by $\mathbf{4.52}$ and $\mathbf{13.44}$, respectively. Additionally, our FSC also achieves promising results. Equipping both approaches can further mitigate performance degradation.
@@ V-D1 Effect of TM
For a more detailed analysis, Fig. 11 (Left) reveals that our methods induce significantly less temporal feature disturbance compared to PTQD. This underscores the efficacy of our motivation and contribution to maintaining temporal features. Further insights are gained by examining the cosine similarity between outputs from the $i$-th Residual Bottleneck Blocks pre- and post-quantization. Comparing our methods with PTQD, Fig. 11 (Right) illustrates that our approaches result in significantly lower output errors in the Residual Bottleneck Block. It is also essential to note that these points involve the accumulated errors from multiple denoising iterations in diffusion models. Therefore, our methods are able to significantly eliminate the cumulative error and quantization error by mitigating temporal feature disturbance within the corresponding block for performance enhancements.
V-D2 Effect of Calibration Methods for TM
Given the variety of available methods to determine the optimal step size for activations during calibration, we explore several approaches and evaluate both their performance and the GPU time consumed. As detailed in Tab. VII, the performance differences between Min-Max and the best-performing ones are minimal, e.g., a mere $\mathbf{0.06}$ FID gap under W4444A8888. Therefore, we opt for the simplest and most efficient Min-Max [61] as our specific calibration strategy, striking a balance between calibration time and effectiveness. Notably, we have found that PTQD and Q-Diffusion cost $4.68$ and $5.29$ GPU hours in their PTQ methods under W4444A8888 quantization, respectively. However, our TIB-based maintenance only spends $\mathbf{2.32}$ GPU hours ($\mathbf{2\times}$ speedup) due to our efficient calibration and reconstruction without extra time consumption compared with previous approaches.
TABLE VII: Different range estimation methods for calibration of TM. We report the GPU hours consumed during calibration in the table. Weight reconstruction with our TIAR ($\mathbf{2.20}$ GPU hours) is used before calibration with FSC. The indices represent the improvements between the best-performing methods and Min-Max [61].
@@ V-D2 Effect of Calibration Methods for TM
V-D3 Effect of Calibration Methods for CM
We also investigate different calibration methods for our Cache-based Maintenance. From Tab. VIII, we observe that LSQ outperforms the others, achieving a $\mathbf{13.3\%}$ increase in SQNR compared to MSE under W8888A8888 settings. Hence, we choose LSQ for calibration. Note that each pre-computed temporal feature can be processed in parallel, enabling the algorithm to achieve rapid runtimes of $\mathbf{<0.3}$ hours on a single $H800$ $80$G GPU.
@@ V-D3 Effect of Calibration Methods for CM
V-D4 Effect of Loss Functions for DS
Since the loss function $\mathcal{L}_{t,i}^{temporal}$ in Eq. (15) measures the error of the temporal feature, we assess various loss functions to enhance performance. As shown in Tab. IX, different loss functions produce comparable results, with less than $\mathbf{1\%}$ FID and sFID discrepancies among them. Moreover, all results in the table outperform either TIB-based or Cache-based Maintenance, demonstrating the robustness and effectiveness of the selection strategy.
TABLE IX: Effect of different loss functions in Eq. (15). We employ MSE [103], aligning with Eq. (14) in our framework.
@@ V-D4 Effect of Loss Functions for DS
V-D5 Effect of Selection Proportion
Beginning with Cache-based Maintenance, we select a proportion of the temporal feature with the lowest $\tau_{t,i}$ for TIB-based Maintenance. Then we increase the proportion progressively until the selection pattern is equivalent to TIB-based Maintenance. As shown in Fig. 12, the lowest FID and highest SQNR were achieved by DS compared with other points. Furthermore, selecting maintenance strategies that minimize temporal feature errors can help enhance performance. This investigation also validates that temporal feature maintenance is crucial.
Figure 12: Performance from CM $\rightarrow$ DS $\rightarrow$ TM. The x-axis represents the proportion of temporal features applied using TM.
V-D6 Generation with Advanced Samplers
Apart from using the DDIM sampler [81], we also utilize a variant of DDPM [22] called PLMS [51] on the CelebA-HQ $256\times 256$ dataset [32]. This better demonstrates the superiority of our framework compared to previous works. From Tab. X, the introduced DS substantially reduces FID and sFID, surpassing PTQD by margins of $\mathbf{12.61}$ and $\mathbf{7.08}$, respectively. Furthermore, we present experiments performed with the DPM++ solver [54]. As shown in Tab. XI, our framework consistently outperforms existing methods.
@@ V-D6 Generation with Advanced Samplers
VI Conclusions & Limitations
@@
In terms of limitations, we have discovered that other semantic features introduced in conditional diffusion models possess physical significance and impact the generation effect. Nevertheless, these features are often overlooked in current methodologies. We plan to address this in future research. Furthermore, while temporal feature maintenance is effective in both PTQ and QAT scenarios, our current study focuses primarily on the PTQ setting. Future efforts will extend to the QAT setting to pursue lower-bit quantization and further performance improvements.
Appendix
We organize the supplementary material as follows:
  • In Sec. A, we provide a detailed relationship between temporal feature errors and final generation quality;
  • In Sec. B, we conduct more experiments to validate the inappropriate reconstruction target of previous methods;
  • In Sec. C, we explore the effectiveness of our framework on video generation;
  • In Sec. D, we provide a guideline on when to use which proposed strategies;
  • In Sec. E, we conduct experiments to show the superiority of our framework compared to previous PTQ methods under sub-$4$-bit quantization;
  • In Sec. F, we apply our framework to the advanced step-distilled model (i.e., FLUX.1-Schnell);
  • In Sec. G, we deploy our quantized models on edge and mobile devices;
@@ A Relationship between Temporal Feature Errors and Final Generation Quality
In this section, we sort out the relationship between temporal feature error and final generation quality. First, we define temporal feature disturbance as a significant temporal feature error (Sec. IV-A). Then we show that temporal feature disturbance (i.e., significant temporal feature error) accounts for temporal information mismatch (Sec. IV-B), which leads to trajectory deviation (Sec. IV-B). In Fig. 4, the deviated trajectory can result in noisy images gradually shifting from a normal denoising trajectory. Thus, the image quality degrades severely. Moreover, as reflected in Fig. 2, with temporal feature error increasing [^11] Obviously, temporal feature error which can also be defined by $\mathcal{L}^{temporal}_{t,i}$ in Eq. (14) increases with the increase of noise level $\lambda$., generation quality measured by FID score deteriorates more sharply.
@@ B Inappropriate Reconstruction Target Validation
We further conduct the validation experiments for the baselines (i.e., PTQ4DM [78], Q-Diffusion [45], and PTQD [19]) in Tab. II. As shown in Tab. I, the Freeze strategy still significantly improves the performance, akin to that in Tab. I. For example, it decreases FID and sFID by $\mathbf{1.18}$ and $\mathbf{5.92}$ for W4444A8888 Q-Diffusion. Hence, the results further validate our analysis and make it more convincing.
@@ C Exploration for Video Generation
Advanced Video diffusion models employ DiT [68] as their backbones. We have found that this type of backbone has temporal features that can be formulated by Eq. (7), where $g_{i}(\cdot)$ is a simple multi-layer perception (MLP) network. Moreover, temporal features and modules generating them are also independent of any $\mathbf{x}_{t}$. Thus, we can directly use our framework for these video diffusion models. Due to similar architectures and algorithms, we believe our approach can also achieve superior performance. As shown in Tab. II, we conduct experiments on OpenSora [102] with $100$-steps DDIM [81] and cfg scale of $4.0$. We select representative metrics and measure them on $10$ examples in the OpenSora prompts set [102], the same as the previous study [99]. Specifically, following EvalCrafter [52], we select CLIPSIM and CLIP-Temp to measure the text-video alignment and temporal semantic consistency. DOVER [91]’s video quality assessment (VQA) metrics to evaluate the generation quality from aesthetic and technical aspects are also involved. $\Delta$Flow Score [52] is used for evaluating the temporal consistency. We employ $8$ prompts from the OpenSora prompt set to quantize the model. Other quantization settings are the same as those of PTQ4DiT [92]. Compared with methods tailored for DiT architecture in Tab. II, our framework outperforms them across all metrics. We will explore more about video generation in the future.
@@ D Guidelines on When to Use Which Proposed Strategies
We provide a guideline here on when to use which maintenance strategy. First, we clarify that the time overhead  [^12] The additional time overhead only related to the process to obtain quantized temporal features is unrelated to $T$. of the two maintenance strategies is all less than $\mathbf{5\%}$ as shown in Fig. 9. Only considering memory overhead, both methods might be suitable for different scenarios. For example, to accommodate a dynamic timestep range of $1\sim 1000$, solely storing temporal features without $\{g_{i}\}_{i=1,\ldots,n}\cup\{h\}$ (i.e., Cache-based Maintenance) for W4444A8888 LDM-8 on LSUN-Churches $256\times 256$ [94] requires an additional $\mathbf{20.7}$MB. It is a significant extra cost amounting to roughly $\mathbf{14.7\%}$ of the quantized UNet size (i.e., $140.9$MB). In contrast, TIB-based Maintenance barely incurs less than $\mathbf{1\%}$ memory overhead in this scenario. However, Cache-based Maintenance can also require less storage in other cases, e.g., $\mathbf{26.65}$MB memory reduction for Stable Diffusion than TIB-based Maintenance as depicted in Fig. 10. In conclusion, users can choose these two maintenance methods for deployment according to their specific hardware, diffusion backbone, inference arguments, etc. Without considering any resource restriction, we suggest users use both methods with our selection approach for higher performance.
@@ E Sub-4-bit Quantization Results
TABLE III: Sub-$4$-bit quantization results for unconditional/class-conditional image generation with LDM. Experimental details are the same as those in the main text.
@@
TM 2/4 38.44 +8.32 162.33 -43.00 174.42 -42.64 64.20 -14.24 29.05 -8.58 50.01 -5.59 35.44 -8.79
CM 2/4 42.37 +12.25 150.21 -55.12 179.04 -38.02 60.24 -18.20 28.41 -9.22 49.46 -6.14 33.90 -10.33
DS 2/4 46.02 +15.90 147.32 -58.01 170.11 -46.95 56.38 -22.06 27.24 -10.39 47.42 -8.18 31.04 -13.19
For sub-$4$-bit settings, we employ W4444A4444 and W2222A4444 quantization with LDM on ImageNet $256\times 256$ [9], LSUN-Bedrooms $256\times 256$ [9], and LSUN-Churches $256\times 256$ [94]. The results in Tab. III show the significant performance improvement achieved by our method compared with other PTQ baselines. Specifically, the superiority of our method to other baselines widens when the bit-width decreases. However, we have to clarify that most current hardware does not support sub-$4$-bit quantization. Moreover, in these extremely low-bit settings, QAT, which also re-trains or fine-tunes model weights, is a more common choice. As discussed in the conclusion section, we believe the insight of temporal feature maintenance can also be generalized to QAT settings for further improving sub-$4$-bit quantization performance. This can be left as our future work.
@@
Methods #Bits (W/A) FLUX.1-Schnell ($\mathbf{T=1}$) FLUX.1-Schnell ($\mathbf{T=4}$)
@@
As shown in Tab. V, our method for SD-XL-turbo [77] with a single sampling step distilled from SD-XL-base [70] shows satisfactory boosts. The improvement is even $\mathbf{>2}$ points larger than that for the $50$-step SD-XL-base under the W4444A8888 setting. Here, we further conduct experiments for FLUX.1-Schnell [43] in this section. Specifically, we employ $512$ COCO prompts for quantization. Other settings are the same as those of SD-XL-turbo [77]. The results in Tab. IV further validate the essential role of temporal features in this kind of scenario.